Having trouble with this one
Suppose you deposit $20,000 for 3 years at a rate of 8%.
If a bank compounds continuous, then the formula becomes simpler, that is A = Pe^rt where e is a constant and equals approximately 2.7183
Now suppose, instead of knowing t, we know that the bank returned to us $25,000 with the bank compounding continuously. Using logarithms, find how long we left the money in the bank (find t).
Answer:
Show work in this space

Answer by fractalier(2101)

Okay, from the equation A = Pe^rt, let's plug in what we know. Thus
25000 = 20000(e^(.08t))
First divide by 20000
1.25 = e^(.08t)
Now take the ln of both sides
ln 1.25 = .08t
Now divide by .08 and calculate
t = (ln 1.25) / .08
t = 2.79 years